OpenStudy (anonymous):

the expression 6s^2 represents the surface area of a cube with edges of length s. what is the surface area of a cube with each edge length 8.)3 inches 9.)1.5 meters

OpenStudy (bahrom7893):

Just substitute.

OpenStudy (bahrom7893):

They're asking you for: 6*3^2 and 6*1.5^2

OpenStudy (anonymous):

ya?

OpenStudy (bahrom7893):

Yes.

OpenStudy (anonymous):

for 8 i got 6s^2=6(3)=18^2 for 8.) I got answer 324 ft is that right???

OpenStudy (bahrom7893):

6s^2 = 6*(3)^2=6*9=54 Do the next one yourself.

OpenStudy (anonymous):

waz the label?

OpenStudy (bahrom7893):

What label?

OpenStudy (anonymous):

on the 54 answer waz thelabel after 54?

OpenStudy (bahrom7893):

8) 54. Is that what you're asking for?

OpenStudy (anonymous):

yeah 54 what? 54 inches 54 feet 54 cm ????

OpenStudy (bahrom7893):

inches^2

OpenStudy (anonymous):

i got 13.5in^2

OpenStudy (anonymous):

for 9. is that right??